# Two-Time-Scale Learning Dynamics: A Population View of Neural Network Training

### 20 Mar 2026 v1

@@ Abstract
Population-based learning paradigms, including evolutionary strategies, Population-Based Training (PBT), and recent model-merging methods, combine fast within-model optimisation with slower population-level adaptation. Despite their empirical success, a general mathematical description of the resulting collective training dynamics remains incomplete. We introduce a theoretical framework for neural network training based on two-time-scale population dynamics. We model a population of neural networks as an interacting agent system in which network parameters evolve through fast noisy gradient updates of SGD/Langevin type, while hyperparameters evolve through slower selection–mutation dynamics. We prove the large-population limit for the joint distribution of parameters and hyperparameters and, under strong time-scale separation, derive a selection–mutation equation for the hyperparameter density. For each fixed hyperparameter, the fast parameter dynamics relaxes to a Boltzmann–Gibbs measure, inducing an effective fitness for the slow evolution. The averaged dynamics connects population-based learning with bilevel optimisation and classical replicator–mutator models, yields conditions under which the population mean moves toward the fittest hyperparameter, and clarifies the role of noise and diversity in balancing optimisation and exploration. Numerical experiments illustrate both the large-population regime and the reduced two-time-scale dynamics, and indicate that access to the effective fitness, either in closed form or through population-level estimation, can improve population-level updates.
@@ 1 Introduction
Figure 1: Two-scale modeling pipeline. Large-population limits ($N\to\infty$) connect the agent-based neural-network dynamics to PDE descriptions; a subsequent two-scale separation yields reduced hyperparameter dynamics and an averaged PDE.
@@
From a theoretical point of view, minibatch stochastic gradient descent admits a well-known interpretation as a discretised Langevin diffusion (Mandt et al., 2017; Li et al., 2019b), whose stationary distribution is asymptotically Gibbsian (Chaudhari and Soatto, 2018). This observation provides a natural foundation for two-time-scale modelling approaches, in which fast stochastic gradient dynamics rapidly equilibrate model parameters, while slower population-level mechanisms act on hyperparameters or higher-level traits. At a macroscopic level, such dynamics are closely related to classical selection–mutation models and population transport equations (Ackleh et al., 2016; Alfaro and Veruete, 2019; Bencheikh and Jourdain, 2022; Del Moral and Miclo, 2001; Del Moral, 2001; Perthame, 2007). This separation enables a mathematically tractable description, clarifies the role of diversity and stability, and connects evolutionary heuristics to replicator–mutation dynamics and transport equations.
@@
  • Large-population limit (Theorem 4). We introduce a continuous-time interacting agent model for PBT in which network parameters evolve via Langevin/SGD dynamics and hyperparameters evolve via selection–mutation. We rigorously prove that the multi-agent system converges to a kinetic PDE description as the population size $N\to\infty$ via coupling techniques and a truncated Wasserstein distance.
  • Two-time-scale reduction (Section 3). Under strong time-scale separation ($\varepsilon\to 0$), we derive a closed selection–mutation PDE for the hyperparameter density $\rho(t,h)$, driven by an effective fitness $\overline{\mathcal{F}}(h)$ that averages the fitness against the Boltzmann–Gibbs equilibrium of the fast parameter dynamics. We show this reduction connects PBT to bilevel optimisation (Lemma 8) and to classical replicator–mutator equations from mathematical biology (Section 3.2).
@@
The remainder of the paper is organised as follows. In Section 2, we introduce the stochastic particle model describing a population of neural networks undergoing both training and evolutionary dynamics. We also derive the corresponding kinetic model in the many-agent limit $N\to\infty$, first formally and then rigorously. In Section 3, we analyse the two-time-scale limit, leading to a reduced macroscopic selection–mutation dynamics for hyperparameters and its connection with bilevel optimisation and replicator–mutator models. In Section 4, we study this reduced dynamics and identify assumptions under which the population mean moves toward the fittest hyperparameter configuration. Numerical examples are presented in Section 5, and concluding remarks and perspectives are discussed in Section 6.
2 Mathematical modelling of population-based training
@@ 2.2 Description in the many-agent limit
A PDE model.
As discussed earlier, the parameter $\tau\in(0,1]$ describes how frequently each agent has the possibility of undergoing a jump. It can be interpreted as a time-discretisation parameter for the collective evolutionary dynamics where, in the limit $\tau\to 0$, jumps may occur at any time and asynchronously. As $N\to\infty$ and $\tau\to 0$, we formally obtain a continuum description of Algorithm 1, in which the time evolution of the kinetic distribution $f$ is determined by both the training and evolutionary operators
@@ 2.3 Propagation of chaos and convergence
We show now that the connection between Algorithm 1 and the kinetic model (6) can be made rigorous. We will consider weak measure solutions to (6), see Definition 12 for a precise characterisation. We restrict to analysis to the case of SGD parameters training with isotropic noise.
@@
To compare the empirical distribution $f^{N,\tau}$ (we stress now also the dependence on $\tau\in(0,1)$) and the kinetic distribution $f$, we will employ the Bounded-Lipschitz (BL) norm between measures, also known as flat, or Dudley metric (Dudley, 2018). For any Borel measure $\mu\in\mathcal{M}(\mathbb{R}^{d})$ it is defined as
@@ 2.4 Modelling extensions
Substitution of NNs performing poorly.
@@
Rank-based selection.
@@
Therefore, at least formally, it is possible to derive kinetic description of the collective training dynamics where rank-based selection is used, as done in Section 2.2 for fitness-based selection. A rigorous convergence analysis is more delicate, and we leave it for future work. In the literature, rank-based particle interactions have been considered in (Bencheikh and Jourdain, 2022; Demircigil and Tomasevic, 2025).
@@ 3 Two-time-scale limit
$\int e^{\mathcal{F}}\,df=\iint e^{\mathcal{F}(\theta,h)}\,\rho(t,h)\,\mu^{\infty}_{\mathcal{L}}(\theta|h)\,d\theta\,dh=\int\exp\!\left(\log\int e^{\mathcal{F}(\theta,h)}\mu^{\infty}_{\mathcal{L}}(\theta|h)\,d\theta\right)\rho(t,h)\,dh=\int e^{\overline{\mathcal{F}}}\,d\rho\,$,
@@ 3.2 Diffusion–reaction approximation in the slow mutation regime
which has been largely studied in the context of evolutionary equations in mathematical biology (Perthame, 2007). The mutation mechanism is now described by a Laplacian term, while selection is described by a source term where particles at $h$ are created with rate $\overline{\mathcal{F}}(h)$ destroyed uniformly with a uniform rate $\langle\overline{\mathcal{F}},\rho\rangle$. Note that the terms are balanced so that the total mass remains constant during the evolution. Equations of type (16) are known in the sampling literature also as birth-death or replicator dynamics (Lu et al., 2023), and can be cast as gradient flows with respect to the Hellinger–Kantorovich (Wasserstein–Fisher–Rao) geometry when additional drift terms are present (Liero et al., 2018; Gallouët et al., 2019). The slow mutation limit therefore reveals a remarkable connection between genetic-type algorithms and gradient flows.
@@ 5.1 Illustrative examples
Figure 3 illustrates the many-agent limit of the PBT dynamics for $N=10^{3},10^{4},10^{5}$. As predicted by Theorem 4, the evolution becomes less noisy as $N$ increases, suggesting convergence toward the deterministic kinetic model (6). Since the mutation strength does not decrease during the computation, the large-time distribution of the hyperparameters is approximately Gaussian and centered around the optimal values $h_{0}=0$ and $h_{1}=0$. This is consistent with the convergence analysis of Theorem 10.
@@ B.2 Limit $N\to\infty$
Non-linear particle system.
Initial data is the same: $\mathbf{X}_{0}=\overline{\mathbf{X}}_{0}$. Then, as before, the non-linear system evolves in the time interval $(t_{n},t_{n+1})$ as
@@
Grönwall-type argument.
We now consider the empirical measure $\overline{f}^{N,\tau}(t)$ associated with the non linear system $\overline{\textbf{X}}_{t}$, and track its distance from $f^{N,\tau}(t)$. Recall that since $\overline{f}^{N,\tau}(t)$ is made of independent i.i.d particles with law $f^{\tau}(t)$, and so $\|\overline{f}^{N\tau}(t)-f^{\tau}(t)\|_{\mathrm{BL}}\lesssim\epsilon(N)$ by Lemma 13.
@@
$\mathbb{E}\left(|X_{t_{n+1}}^{i}-\overline{X}_{t_{n+1}}^{i}|\wedge 1\right)$ $=(1-\tau)\mathbb{E}\left(|X_{t_{n+1}^{-}}^{i}-\overline{X}_{t_{n+1}^{-}}^{i}|\wedge 1\right)+\tau\mathbb{E}\left(|X_{n}^{*}\left(\mathbf{w}_{n},\mathbf{X}_{t_{n+1}^{-}},\alpha\right)-X^{\mathbf{j}(\mathbf{w}_{m},\alpha_{n}^{i})}|\wedge 1\right)$
$=(1-\tau)\mathbb{E}\left(|X_{t_{n+1}^{-}}^{i}-\overline{X}_{t_{n+1}^{-}}^{i}|\wedge 1\right)+\tau\left\|G_{\mathcal{F}}\left[f^{N,\tau}(t_{n+1}^{-})\right]-G_{\mathcal{F}}\left[f^{\tau}(t_{n+1}^{-})\right]\right\|_{\mathrm{BL}}\,$
@@ C.1 Proof of Theorem 10
Evolution of second moments.
We denote the energy, or second moment, of $\rho$ as $\textup{En}(\rho)=M_{2}(\rho)=\int|h|^{2}\rho(dh)$. Thanks to Assumption 9 and (21), we have
@@
Evolution of mass over $B(h^{\star},r)$.
Thanks to (35), we can apply Lemma 16 to the PDE solution $\rho(t)$, and obtain that for any $r>0$ there exists $\delta>0$ such that
@@
Quantitative Laplace principle.
In the following, we denote the weighted mean as $\textup{m}^{\alpha}[\rho]:=\textup{m}\left(G_{\alpha\overline{\mathcal{F}}}[\rho(t)]\right)$ for simplicity. Let $tol>0$ be fixed. By the quantitative Laplace principle (22), for any $\rho\in\mathcal{P}_{2}(\mathbb{R}^{d_{h}})$, under the inverse continuity Assumption 9 (iii) we have that there exists $r_{tol}>0$ and $q_{tol}>0$ sufficiently small such that

### 25 Mar 2026 v2

@@ Abstract
Population-based learning paradigms, including evolutionary strategies, Population-Based Training (PBT), and recent model-merging methods, combine fast within-model optimisation with slower population-level adaptation. Despite their empirical success, a general mathematical description of the collective dynamics remains incomplete. We propose a theoretical framework for neural network training based on two-time-scale multi-agent dynamics, where parameters follow fast noisy SGD/Langevin updates and hyperparameters evolve via slower selection–mutation dynamics. We prove the large-population limit for the joint distribution of parameters and hyperparameters and, under strong time-scale separation, derive a selection–mutation equation for the hyperparameter density. For each fixed hyperparameter, the fast parameter dynamics relaxes to a Boltzmann–Gibbs measure, inducing an effective fitness for the slow evolution. The averaged dynamics connects population-based learning with bilevel optimisation and classical replicator–mutator models, yields conditions under which the population mean moves toward the fittest hyperparameter, and clarifies the role of noise and diversity in balancing optimisation and exploration. Numerical experiments illustrate both the large-population regime and the reduced two-time-scale dynamics, and indicate that access to the effective fitness, either in closed form or through population-level estimation, can improve population-level updates.
@@ 1 Introduction
Figure 1: Two-scale modelling pipeline. Large-population limits ($N\to\infty$) connect the agent-based neural-network dynamics to a two-scale kinetic PDE for the joint density $f(t,\theta,h)$ over parameters and hyperparameters, in which a training operator and a selection–mutation evolutionary operator act on different time scales. A two-scale separation ($\varepsilon\to 0$) averages out the fast parameter dynamics, which relaxes to a Boltzmann–Gibbs equilibrium $\mu^{\infty}_{\mathcal{L}}(\theta|h)$, and yields a closed averaged PDE for the hyperparameter density $\rho(t,h)$ driven by the effective fitness $\overline{\mathcal{F}}(h)$.
@@
From a theoretical point of view, minibatch stochastic gradient descent admits a well-known interpretation as a discretised Langevin diffusion (Mandt et al., 2017; Li et al., 2019b; Chizat, 2022), whose stationary distribution is asymptotically Gibbsian (Chaudhari and Soatto, 2018). This observation provides a natural foundation for two-time-scale modelling approaches, in which fast stochastic gradient dynamics rapidly equilibrate model parameters, while slower population-level mechanisms act on hyperparameters or higher-level traits. At a macroscopic level, such dynamics are closely related to classical selection–mutation models and population transport equations (Ackleh et al., 2016; Alfaro and Veruete, 2019; Bencheikh and Jourdain, 2022; Del Moral and Miclo, 2001; Del Moral, 2001; Perthame, 2007). This separation enables a mathematically tractable description, clarifies the role of diversity and stability, and connects evolutionary heuristics to replicator–mutation dynamics and transport equations.
@@
  • Large-population limit (Theorem 4). We introduce a continuous-time interacting agent model for PBT in which network parameters evolve via Langevin/SGD dynamics and hyperparameters evolve via selection–mutation. We rigorously prove that the multi-agent system converges to a two-scale kinetic PDE description as the population size $N\to\infty$ via coupling techniques and a truncated Wasserstein distance.
  • Two-time-scale reduction (Section 3). Under strong time-scale separation ($\varepsilon\to 0$), we derive an averaged selection–mutation PDE for the hyperparameter density $\rho(t,h)$, driven by an effective fitness $\overline{\mathcal{F}}(h)$ that averages the fitness against the Boltzmann–Gibbs equilibrium of the fast parameter dynamics. We show this reduction connects PBT to bilevel optimisation (Lemma 8) and to classical replicator–mutator equations from mathematical biology (Section 3.2).
@@
The remainder of the paper is organised as follows. In Section 2, we introduce the stochastic particle model describing a population of neural networks undergoing both training and evolutionary dynamics. We also derive the corresponding two-scale PDE model in the many-agent limit $N\to\infty$, first formally and then rigorously. In Section 3, we analyse the two-time-scale limit, leading to a reduced macroscopic selection–mutation dynamics for hyperparameters and its connection with bilevel optimisation and replicator–mutator models. In Section 4, we study this reduced dynamics and identify assumptions under which the population mean moves toward the fittest hyperparameter configuration. Numerical examples are presented in Section 5, and concluding remarks and perspectives are discussed in Section 6.
2 Modelling population-based training
@@ 2.2 Description in the many-agent limit
As discussed earlier, the parameter $\tau\in(0,1]$ describes how frequently each agent has the possibility of undergoing a jump. It can be interpreted as a time-discretisation parameter for the collective evolutionary dynamics where, in the limit $\tau\to 0$, jumps may occur at any time and asynchronously. As $N\to\infty$ and $\tau\to 0$, we formally obtain a continuum description of Algorithm 1, in which the time evolution of the distribution $f$ is determined by both the training and evolutionary operators
@@
Equations of this type are inspired by recent applications of kinetic theory (Pareschi and Toscani, 2013) and are closer in spirit to kinetic formulations of simulated annealing (Pareschi, 2024; Chizat, 2022) and genetic algorithms (Borghi and Pareschi, 2025).
@@ 2.3 Propagation of chaos and convergence
We show now that the connection between Algorithm 1 and the two-scale PDE model (6) can be made rigorous. We will consider weak measure solutions to (6), see Definition 12 for a precise characterisation. We restrict to analysis to the case of SGD parameters training with isotropic noise.
@@
To compare the empirical distribution $f^{N,\tau}$ (we stress now also the dependence on $\tau\in(0,1)$) and the distribution $f$, we will employ the Bounded-Lipschitz (BL) norm between measures, also known as flat, or Dudley metric (Dudley, 2018). For any Borel measure $\mu\in\mathcal{M}(\mathbb{R}^{d})$ it is defined as
@@ 2.4 Modelling extensions
2.4.1 Substitution of NNs performing poorly
@@
2.4.2 Rank-based selection
@@
Therefore, at least formally, it is possible to derive a similar two-scale PDE description of the collective training dynamics where rank-based selection is used, as done in Section 2.2 for fitness-based selection. A rigorous convergence analysis is more delicate, and we leave it for future work. In the literature, rank-based particle interactions have been considered in (Bencheikh and Jourdain, 2022; Demircigil and Tomasevic, 2025).
@@ 3 Two-time-scale limit
$\int e^{\mathcal{F}}\,df$ $=\iint e^{\mathcal{F}(\theta,h)}\,\rho(t,h)\,\mu^{\infty}_{\mathcal{L}}(\theta|h)\,d\theta\,dh$
$=\int\exp\!\left(\log\int e^{\mathcal{F}(\theta,h)}\mu^{\infty}_{\mathcal{L}}(\theta|h)\,d\theta\right)\rho(t,h)\,dh=\int e^{\overline{\mathcal{F}}}\,d\rho\,$,
@@ 3.2 Diffusion–reaction approximation in the slow mutation regime
which has been largely studied in the context of evolutionary equations in mathematical biology (Perthame, 2007). The mutation mechanism is now described by a Laplacian term, while selection is described by a source term where particles at $h$ are created with rate $\overline{\mathcal{F}}(h)$ destroyed uniformly with a uniform rate $\langle\overline{\mathcal{F}},\rho\rangle$. Note that the terms are balanced so that the total mass remains constant during the evolution. Equations of type (16) are known in the sampling literature also as birth-death or replicator-mutation dynamics (Lu et al., 2023), and can be cast as gradient flows with respect to the Hellinger–Kantorovich (Wasserstein–Fisher–Rao) geometry when additional drift terms are present (Liero et al., 2018; Gallouët et al., 2019). The slow mutation limit therefore reveals a remarkable connection between genetic-type algorithms and gradient flows.
@@ 5.1.1 Quadratic function
Figure 3 illustrates the many-agent limit of the PBT dynamics for $N=10^{3},10^{4},10^{5}$. As predicted by Theorem 4, the evolution becomes less noisy as $N$ increases, suggesting convergence toward the deterministic kinetic PDE model (6). Since the mutation strength does not decrease during the computation, the large-time distribution of the hyperparameters is approximately Gaussian and centered around the optimal values $h_{0}=0$ and $h_{1}=0$. This is consistent with the convergence analysis of Theorem 10.
@@ B.2 Limit $N\to\infty$
Non-linear particle system. Initial data is the same: $\mathbf{X}_{0}=\overline{\mathbf{X}}_{0}$. Then, as before, the non-linear system evolves in the time interval $(t_{n},t_{n+1})$ as
@@
Grönwall-type argument. We now consider the empirical measure $\overline{f}^{N,\tau}(t)$ associated with the non linear system $\overline{\textbf{X}}_{t}$, and track its distance from $f^{N,\tau}(t)$. Recall that since $\overline{f}^{N,\tau}(t)$ is made of independent i.i.d particles with law $f^{\tau}(t)$, and so $\|\overline{f}^{N\tau}(t)-f^{\tau}(t)\|_{\mathrm{BL}}\lesssim\epsilon(N)$ by Lemma 13.
@@
$\mathbb{E}\left(|X_{t_{n+1}}^{i}-\overline{X}_{t_{n+1}}^{i}|\wedge 1\right)$ $=(1-\tau)\mathbb{E}\left(|X_{t_{n+1}^{-}}^{i}-\overline{X}_{t_{n+1}^{-}}^{i}|\wedge 1\right)$
$\qquad+\tau\mathbb{E}\left(|X_{n}^{*}\left(\mathbf{w}_{n},\mathbf{X}_{t_{n+1}^{-}},\alpha\right)-X^{\mathbf{j}(\mathbf{w}_{m},\alpha_{n}^{i})}|\wedge 1\right)$
$=(1-\tau)\mathbb{E}\left(|X_{t_{n+1}^{-}}^{i}-\overline{X}_{t_{n+1}^{-}}^{i}|\wedge 1\right)$
$\qquad+\tau\left\|G_{\mathcal{F}}\left[f^{N,\tau}(t_{n+1}^{-})\right]-G_{\mathcal{F}}\left[f^{\tau}(t_{n+1}^{-})\right]\right\|_{\mathrm{BL}}\,$
@@ C.1 Proof of Theorem 10
Evolution of second moments. We denote the energy, or second moment, of $\rho$ as $\textup{En}(\rho)=M_{2}(\rho)=\int|h|^{2}\rho(dh)$. Thanks to Assumption 9 and (21), we have
@@
Evolution of mass over $B(h^{\star},r)$. Thanks to (35), we can apply Lemma 16 to the PDE solution $\rho(t)$, and obtain that for any $r>0$ there exists $\delta>0$ such that
@@
Quantitative Laplace principle. In the following, we denote the weighted mean as $\textup{m}^{\alpha}[\rho]:=\textup{m}\left(G_{\alpha\overline{\mathcal{F}}}[\rho(t)]\right)$ for simplicity. Let $tol>0$ be fixed. By the quantitative Laplace principle (22), for any $\rho\in\mathcal{P}_{2}(\mathbb{R}^{d_{h}})$, under the inverse continuity Assumption 9 (iii) we have that there exists $r_{tol}>0$ and $q_{tol}>0$ sufficiently small such that